OpenStudy (anonymous):

what is an equation of the line, in point slope form, that passes through the given point and has the give slope? point: (7,3); slope: 4/7 y-3=4/7(x-7) y-7=4/7(x-3) y-3=4/7(x+7) y-3=-4/7(x-7)

OpenStudy (marigirl):

y-y1=m(x-x1) y-3=(4/7)(x-7)

OpenStudy (marigirl):

where x1 and y1 refer to the coordiantes provided to us and m is the gradient (slope)

OpenStudy (anonymous):

that's right...y-y1=m(x-x1)...... see I can never find the stuff that I did 3 months ago in my book. its ridiculous. thank you for helping :)

OpenStudy (marigirl):

no worries :)